### v3

# SAM 2++: Tracking Anything at Any Granularity

Abstract
Video tracking aims at finding the specific target in subsequent frames given its initial state. Due to the varying granularity of target states across different tasks, most existing trackers are tailored to a single task and heavily rely on custom-designed modules within the individual task, which limits their generalization and leads to redundancy in both model design and parameters. To unify video tracking tasks, we present SAM 2++, a unified model towards tracking at any granularity, including masks, boxes, and points. First, to extend target granularity, we design task-specific prompts to encode various task inputs into general prompt embeddings, and a unified decoder to unify diverse task results into a unified form pre-output. Next, to satisfy memory matching, the core operation of tracking, we introduce a task-adaptive memory mechanism that unifies memory across different granularities. Finally, we introduce a customized data engine to support tracking training at any granularity, producing a large and diverse video tracking dataset with rich annotations at three granularities, termed Tracking-Any-Granularity, which represents a comprehensive resource for training and benchmarking on unified tracking. Comprehensive experiments on multiple benchmarks confirm that SAM 2++ sets a new state of the art across diverse tracking tasks at different granularities, establishing a unified and robust tracking framework.
  ‡Equal contribution.  †Corresponding author (lmwang@nju.edu.cn).
1 Introduction
Video tracking has been a fundamental task in computer vision for decades, aiming to estimate the state of an arbitrary target in video sequences given its initial status. Despite sharing this core objective, the tracking domain has fragmented into several independent sub-tasks based on different target granularities, including Single Object Tracking [lasot, TrackingNet, got10k] (SOT) with bounding box, Video Object Segmentation [davis17, youtube-vos, LVOS_V1] (VOS) with precise pixel-level mask, and Point Tracking [badja, zheng2023pointodyssey, tapvid] with tiny points. This fragmentation based on state granularity has led most video tracking research to focus on a specific task and propose specialized designs only for that task. While this design trend enhances tracking performance, it limits the generalization ability of tracking models across multiple tasks and results in redundancy in both model design and parameters. To unify tasks, current unified vision models typically share feature extraction backbones while employing task-specific branches [zhu2022uniperceiver], convert those tasks into a seq2seq framework [chen2021pix2seq], or share one appearance model for either propagation or association [wang2021UniTrack, yu2024unifiedtt, Unicorn, UNINEXT, OmniTracker]. However, they choose to provide different interfaces for different tasks, rather than seeking a unified visual representation of tracking targets, and ignore the point tracking task.
Unlike them, we observe that these seemingly disparate tracking paradigms fundamentally differ primarily in their state granularity, while sharing the memory matching strategy: the model encodes the previous state into memory, and matches the current features with the stored memory when a new frame is received. Based on this strategy, we decide to unify target states at three different granularities through a uniform memory representation. Recently, Segment Anything Model 2 [ravi2024sam2], a strong foundational model, has been proposed for high-quality video object segmentation given various prompts. Due to its flexible prompt mechanism and powerful mask tracking capabilities, we extend this model to track arbitrary granularity, termed as SAM 2++.
Figure 1: The overall of SAM 2++, including (a) tracking any granularity task, (b) our unified tracking foundation model, and (c) our Tracking-Any-Granularity dataset collected through our data engine. SAM 2++ is capable of tracking targets at any granularity.
Our work includes a model and a dataset (see Fig. 1). To ensure generalized tracking at different granularities, we start by designing task-specific prompts and a unified decoder. Specifically, we introduce corresponding prompts for different tasks in various granularities to encode various task inputs into general prompt embeddings. As for the diverse task output, our unified decoder, which is extended from the Mask Decoder of SAM 2, unifies diverse task results into a unified form pre-output. Next, we found that a simple full parameter-shared approach for task mixing training leads to performance degradation on all tasks, due to the different memory requirements of different tracking tasks. To address this, we introduce a task-adaptive memory mechanism, which adjusts memory representations in response to the unique requirements of each task. This mechanism not only helps to offset the adverse effects of full parameter sharing on the memory mechanism but also achieves mutual promotion among multiple tasks. Finally, to enable ”tracking granularity” capabilities in video, we utilize a data engine to construct a large and diverse video tracking dataset, termed Tracking-Any-Granularity (TAG). The data engine produces training data through an interactive process, where annotators manually label data at varying intervals in different phases. Subsequently, after training on the datasets at different phases, the model is used to annotate the remaining frames, achieving efficient and accurate expansion of the dataset. Unlike most existing video tracking datasets, our dataset provides high-quality annotations at three granularities, including segmentation masks, bounding boxes, and key points, resulting in a vital resource for training and benchmarking unified tracking models. Extensive experiments on several benchmarks from various tasks demonstrate that SAM 2++ enables tracking targets at any granularity with a unified model architecture and consistently outperforms task-specific models in all three tasks.
The main contributions are summarized as follows:
  • We propose a unified framework, termed SAM 2++, towards tracking targets at any granularity by task-specific prompts, a unified decoder, and a task-adaptive memory mechanism for various granularities.
  • We build a data engine that produces training data through an interactive process, resulting in a new large-scale object tracking dataset, Tracking-Any-Granularity (TAG), with high-quality annotations in various granularities.
  • Experiments show that SAM 2++ enables accurate tracking at various granularities, consistently surpassing the performance of task-specific models.
2 Related work
Segment Anything Model. SAM [kirillov2023seganysam1] is a foundational model for high-quality segmentation given various prompts, and SAM 2 [ravi2024sam2] extends it to video with streaming memory, effectively handling motion and occlusion, and they inspire many variant models In the image domain, HQ-SAM [sam_hq] enhances segmentation quality through a High-Quality Token, SAMRefiner [lin2025samrefiner] improves fine-grained details via a noise-tolerant prompt, CAT-SAM [xiao2024cat] adopts a conditional tuning approach to adapt to specialized image domains, and SAM-Adapter [chen2023samadapter] incorporates lightweight adapters for improved downstream performance. In the video domain, SAM2Long [ding2024sam2long] employs constrained tree search to reduce error accumulation, SAMURAI [yang2024samurai] uses the Kalman filter to select motion-aware memory, while DAM4SAM [videnovic2024distractor] introduces a distractor-aware memory. SAMWISE [cuttano2024samwise] and AL-Ref-SAM-2 [huang2025unleashing] add additional prompts for more referring tasks. Despite these advances, these works remain task-specific, lacking cross-domain generalization and requiring separate implementations for each application.
Unified Vision Models. Recent years have witnessed significant progress in developing unified vision models that handle multiple tasks through shared architectures and demonstrate strong generalizability and flexibility. Pix2Seq [chen2021pix2seq] reformulates vision tasks as sequence generation problems, Uni-Perceiver [zhu2022uniperceiver] establishes unified representation spaces across modalities with shared encoders and decoders. UniTrack [wang2021UniTrack] demonstrates that video tracking tasks can be solved by a single appearance model with task-specific heads, while Unicorn [Unicorn] and UNINEXT [UNINEXT] unify various tracking paradigms through common frameworks with different representations. Despite their impressive capabilities, these unified approaches predominantly focus on object-level tasks while neglecting finer-grained tasks such as point tracking. Furthermore, they do not take into account unifying video tracking tasks with various granularities through a unified visual representation.
3 Preliminaries: Segment Anything Model 2
The Segment Anything Model (SAM) [kirillov2023seganysam1] is a milestone vision foundation model for class-agnostic image segmentation. It flexibly handles various prompts (box, point, mask) by encoding them into a unified embedding and has established an iterative data engine with model-assisted labeling to address dataset limitations. SAM 2 [ravi2024sam2] extends SAM to promptable video segmentation by introducing a streaming memory that stores previous target information and predictions. It comprises four main components: (i) a hierarchical image encoder that encodes each frame ${I}_{img}$ into image embeddings $F_{img}$, (ii) a prompt encoder, (iii) a memory mechanism (memory encoder, memory bank, memory attention), and (iv) a mask decoder for prediction.
Prompt Encoder. SAM 2 follows the prompt encoder design from SAM to support three types of user inputs, including positive/negative points, bounding boxes, and masks. The point prompt ${I}_{point}\in\mathbb{R}^{N_{point}\times 2}$ and box prompt ${I}_{box}\in\mathbb{R}^{2\times 2}$ (seen as two corner points) can be represented as sparse embeddings $\mathcal{P}_{sparse}\in\mathbb{R}^{N_{point}\times C}$} by their point location and learnable embedding parameters $\varepsilon_{sparse}^{point},\varepsilon_{sparse}^{box}$ which encodes the type of each point. As for the mask prompt ${I}_{mask}\in\mathbb{R}^{1\times H\times W}$, the model adopts convolutions to map and downscale them as dense embedding $\mathcal{P}_{dense}\in\mathbb{R}^{C\times H/16\times W/16}$}. In summary, the processing of Prompt Encoder can be written as:
$\mathcal{P}_{sparse}$ $=[\textit{PE}({I}_{point})+\varepsilon_{sparse}^{point};\textit{PE}({I}_{box})+\varepsilon_{sparse}^{box}],$ (1)
$\mathcal{P}_{dense}$ $=\mathbf{Conv}_{dense}({I}_{mask}),$
where the PE represents positional encoding operation.
Mask Decoder. The mask decoder takes prompt embedding $\mathcal{P}_{sparse}$ and $\mathcal{P}_{dense}$, memory-conditioned image embeddings $\bar{F}_{img}\in\mathbb{R}^{C/4\times H/16\times W/16}$ (which we will explain latter), and a set of learnable tokens $\mathcal{E}_{tokens}$ as inputs. The learnable tokens contain an existence token $\varepsilon_{obj}\in\mathbb{R}^{C}$ to predict whether the target exists, an IoU token $\varepsilon_{iou}\in\mathbb{R}^{C}$ to predict the result accuracy, and multiple mask tokens $\varepsilon_{mask}^{N}\in\mathbb{R}^{N\times C}$ used to obtain $N$ mask candidates. To fuse the prompt embedding, a Two-Way Transformer $\mathbf{twTrans}$ [ravi2024sam2] processes them as:
$\tilde{F}_{img},[\tilde{\mathcal{P}}_{sparse};\tilde{\mathcal{E}}_{tokens}]=\mathbf{twTrans}($ (2)
$\bar{F}_{img}+\mathcal{P}_{dense},[\mathcal{P}_{sparse};\mathcal{E}_{tokens}]).$
After that, the output token embeddings $\tilde{\mathcal{E}}_{tokens}$ are split into $\tilde{\varepsilon}_{obj}$ for predicting existence $O_{obj}$, $\tilde{\varepsilon}_{iou}$ for producing IoU scores $O_{IoU}^{N}$, and $\tilde{\varepsilon}_{mask}^{N}$ for generating mask output as:
$M^{i}_{mask}=\text{Interpolate}(\tilde{F}_{img}\cdot\tilde{\varepsilon}_{mask}^{i}),$ (3)
where the $M^{i}_{mask}$ represents the $i_{th}$ candidate mask prediction rated by corresponding iou score.
Memory. The memory encoder $\mathbf{MemEn}$ processes image embedding $F_{img}$ and the mask prediction $M_{mask}^{*}$ with the highest IoU score to generate memory embedding $\bar{\bar{F}}_{img}$ for the processed frame. In addition, it introduces object pointer $\varepsilon_{pointer}\in\mathbb{R}^{C}$, which is transformed from the mask token $\tilde{\varepsilon}_{mask}^{*}$, to provide high-level semantic information. After that, these two kinds of memory are appended to Memory Bank $\mathcal{MB}$ in FIFO mode. To enable the current frame to obtain past target information, the image embeddings ${F}_{img}$ are not directly fed to the Mask Decoder, but instead conditioned on memories from Memory Bank as $\bar{F}_{img}$ by cross-attention in Memory Attention $\mathbf{MemAttn}$.
4 Model
In this section, we present our unified video tracking framework, termed as SAM 2 ++, which extends the SAM 2 model to track any targets in videos at any granularity, including masks, bounding boxes, and points, and the overall pipeline is depicted in Fig. 2. Due to the various task granularities, we introduce task-specific prompts to unify task input in different granularities and the Unified Decoder to unify diverse task results into a unified form pre-output. Next, we found that a fully parameter-shared model training results in performance degradation due to the diverse memory requirements across tasks. To address this, we introduce a task-adaptive memory mechanism that dynamically adjusts memory representations according to each task’s demand, enhancing the multi-task processing capability.
Figure 2: The SAM 2++ architecture. When a new frame is received, the result is conditioned on the new prompt and/or stored memories. The initial target state at any granularity is converted into task-specific prompts for unified input. The Unified decoder predicts the task result for the current frame in unified mask form. Finally, the task-adaptive memory transforms diverse target states into unified memory.
4.1 Unified Task Input and Output Processing
Input Unification via Task-Specific Prompt. Due to the input of the three tracking tasks having inconsistent granularity, we first unify inputs with task-specific prompts for different tasks. The video object segmentation task still adopts mask input ${I}_{mask}$ as its prompt in mask form, and the single object tracking task takes its box input ${I}_{box}$ as the prompt. As for the point tracking task, expect the point coordinates ${I}_{point}$, we add a dense mask ${G}_{point}$ for additional prompt as a Gaussian map centred on the point and parameterised by sigma $\sigma$ and radius $r$ as: $G_{point}=\exp\left(-\frac{\|p-p_{0}\|^{2}}{2\sigma^{2}}\right)\cdot\mathbf{1}_{\{\|p-p_{0}\|\leq r\}}$ to highlights the point in mask form, maintaining consistency with output from Unified Decoder and source for Memory Encoder, which is better than naive {0, 1} mask. More importantly, we gradually decrease the radius and sigma during training to facilitate smoother convergence and more stable learning.
Output Unification via Unified Decoder. To unify the output of various tasks, we extended the Mask Decoder of SAM 2 as Unified Decoder, which also processes memory-conditioned image embeddings, prompt embeddings, and learnable tokens. For the SOT task, the outer box of the mask output $M^{N}_{box}$ cannot be used as task output because the complexity of the mask reduces the accuracy of the box, which focuses on the center point’s position and target scale. Instead, we add a Corner-based Head [yan2021learning_stark], $\mathbf{CornHead}$, which explicitly optimizes for the accuracy and stability of the bounding box and is widely used in the SOT task, to produce box predictions. As for the PT task, rather than direct point coordinates, we obtained point predictions in terms of mask predictions by soft-argmax operation during training or argmax operation during inference. This specific design aims for the output of the point task to be consistent with the source of memory, thereby achieving a unified encoding mask output for memory information of different granularities, and also helps to optimize model training. In summary, our unified decoder can be written as:
$M_{mask}^{i}$ $=\text{Interpolate}(\tilde{F}_{img}\cdot\tilde{\varepsilon}_{mask}^{i}),$ (4)
$B_{box}^{i}$ $=\mathbf{CornHead}(\tilde{F}_{img},\tilde{\varepsilon}_{mask}^{i}),$
$P_{point}^{i}$ $=\text{argmax}(\text{Interpolate}(\tilde{F}_{img}\cdot\tilde{\varepsilon}_{mask}^{i})),$
where $M_{mask}^{i}$, $B_{box}^{i}$, and $P_{point}^{i}$ are task predictions. represent the $i_{th}$ candidate prediction for three tasks, which are rated by their corresponding iou scores $O_{IoU}^{i}$.
4.2 Task-adaptive Memory
Tracking models fundamentally localize targets according to their past states, which requires efficient storage and retrieval ability using a memory-matching paradigm: the model first encodes the previous states into memory, then matches current features with memory to accurately represent the target when processing a new frame. Following this paradigm, our model converts mask outputs into memory with Memory Encoder, then applies cross-attention in Memory Attention to match the current frame feature with the feature stored in the memory bank as:
$\bar{\bar{F}}_{img,\rho}$ $=\mathbf{MemEn}(F_{img},M_{\rho}^{*}),$ (5)
$\mathcal{MB}_{\rho}$ $=\text{FIFO}([\varepsilon_{pointer},\bar{\bar{F}}_{img,\rho}]),$
$\bar{F}_{img}$ $=\mathbf{MemAttn}(\mathcal{F}_{img},\mathcal{MB_{\rho}},\mathcal{MB_{\rho}}),$
where $\rho$ represents different granularities in various tasks. However, the mask outputs $M_{\rho}^{*}$ from the three tasks differ in their requirements: in mask tracking, the mask is a precise segmentation; in box tracking, the mask provides coarse localization to assist the box head; in point tracking, the mask is the Gaussian form of the target point. Based on the above analysis, if the model adopts a full parameter-shared memory module for encoding these diverse mask outputs, it fails to generate task-adaptive memory representations accurately, resulting in memory features failing to meet any requirements of the three tasks.
Figure 3: Annotation pipeline of our Tracking-Any-Granularity dataset.
Figure 4: Examples of Tracking-Any-Granularity Dataset.
Therefore, we propose a task-adaptive memory mechanism, which relaxes the uniformity by decoupling only the memory components: each task has its own convolutional Memory Encoder, and each applies an independent LoRA [DBLP:conf/iclr/HuSWALWWC22/lora] in the transformer-based Memory Attention. This decoupled design effectively meets the diverse needs of different tasks and avoids the performance drop seen in a fully parameter-shared model, while keeping the overall structure consistent, enhancing the multi-task processing capability of the model. Since only a small number of parameters are decoupled, the increase in parameter count is minimal. Notably, experiments show that this design enables multiple tasks to promote each other.
4.3 Training and Inference Details
Training. SAM 2++ performs multi-task training on tracking tasks with different granularity (mask, box, point), initialized from SAM 2 base. We decoupled the memory-related modules, including creating a separate copy of the memory encoder and implementing dedicated LoRA parameters for memory attention for each task. In addition to our Tracking-Any-Granularity dataset, we adopt DAVIS 2017 [davis17], YoutubeVOS 2019 [youtube-vos] and MOSE [MOSE] for the mask task; LaSOT [lasot], GOT10k [got10k], TrackingNet [TrackingNet] and COCO [coco] for the box task; TapVid Kinetics [tapvid], PointOdyssey [zheng2023pointodyssey], and PerceptionTest [patraucean2023perception] for the point task. During training, we use 8-frame sequences with up to 3 targets in the first frame, The first frame and one randomly selected frame serve as conditional frames, receiving either normal prompts or interactive prompts with equal probability. For further details, please refer to the Appendix.
Losses and optimization. For mask tracking, we combine focal and dice losses for mask prediction, L1 loss for IoU prediction, and cross-entropy loss for occlusion prediction. Box tracking extends it with ciou [zheng2020diouciou] and L1 losses for box predictions. Point tracking adds L1 loss for point predictions with soft argmax. In multi-prediction scenarios, we only supervise the task prediction with the lowest combined loss while supervising all IoU predictions. For occluded frames, denoted as $1-\mathbbm{1}_{obj}$, we skip the supervision of the task results and IoU prediction, but maintain occlusion prediction supervision. In summary, the multi-task training loss can be written as:
$\mathcal{L}_{Mask}$ $=(\lambda_{mask}^{focal}\mathcal{L}_{mask}^{focal}+\lambda_{mask}^{dice}\mathcal{L}_{mask}^{dice})\times\mathbbm{1}_{obj}$ (6)
$+\lambda_{IoU}^{L1}\mathcal{L}_{IoU}^{L1}\times\mathbbm{1}_{obj}+\lambda_{obj}^{CE}\mathcal{L}_{obj}^{CE}$
$\mathcal{L}_{Box}$ $=\mathcal{L}_{Mask}+(\lambda_{box}^{ciou}\mathcal{L}_{box}^{ciou}+\lambda_{box}^{L1}\mathcal{L}_{box}^{L1})\times\mathbbm{1}_{obj}$
$\mathcal{L}_{Point}$ $=\mathcal{L}_{Mask}+\lambda_{point}^{L1}\mathcal{L}_{point}^{L1}\times\mathbbm{1}_{obj}$
Inference. During inference, we follow a fully online inference setting where only the ground truth of the first frame serves as the initial prompt without any subsequent corrections and future information. Our model operates on full frames without post-processing strategies like center cropping, which is commonly used in tracking tasks.
5 Data
Figure 5: Statistics on sources and attributes distribution of Tracking-Any-Granularity Dataset. The link in (c) reflects the frequent co-occurrence of multiple attributes in a sequence.
Table 1: Comparison of our datasets with public datasets of three tracking datasets in terms of videos, duration, and annotations.
Dataset Videos Total Len. (Avg) Frames (Avg) Resolution FPS Masks (Avg.) Boxes Points (Avg.) Anno. Method Motivation
[…]
BURST [burst] 2914 1734 (0.60) 624240 (214) $\geq$480p 6 600157 (206) × × Semi-Automatic Multi-Task
LVOS [LVOS_V1] 220 351 (1.60) 126280 (574) 720p 6 156432 (711) × × Manual Long-term
LVOS v2 [LVOS_V2] 720 823 (1.14) 296401 (412) 720p 6 407945 (567) × × Manual Large-scale, long-term
MOSE [MOSE] 2149 443.62 (0.21) $\sim$159600 (73) 1080p 6 431725 (201) × × Semi-Automatic Complext scenarios
YoutubeVOS-19 [youtube-vos] 4453 334.8 (0.08) 120532 (27) 720p 6 197272 (44) × × Manual Large-scale
VOST [vost] 713 252 (0.35) 75547 (106) 1080p 5 175913 (247) × × Semi-Automatic Object transmission
LaSOT [lasot] 1400 1950 (1.39) 3.52M (2506) 720p 30 × 3.52M × Manual Large-scale, long-term
[…]
TrackingNet [TrackingNet] 30643 8400 (0.27) 14431266 (471) 360p 30 × 14431266 × Semi-Automatic Large-scale
UAV123 [uav123] 123 62.5 (0.51) 112578 (915) 720p 30 × 112578 × Semi-Automatic Unmanned aerial vehicles
NfS [NFS] 100 26.58 (0.27) 383K (3830) 720p 240 × 383K × Manual High Frame Rate
OTB-100 [otb] 100 32.8 (0.33) 59040 (590) $\geq$360p 30 × 59040 × Manual Real world
TNL2K [wang2021tnl2k] 2000 691.3 (0.35) 1244340 (622) 720p 30 × 1244340 × Manual Language-based
VastTrack [vasttrack] 50610 11664 (0.23) 4.2M (83) 480p-720p 6 × 4.2M × Manual Abundant categories
[…]
PointOdyssey [zheng2023pointodyssey] 104 120 (1.15) $\sim$216K (2035) 540p 30 × × 49B (0.471B) Automatic Real world, long-term
TAP-Vid Kinetics [tapvid] 1189 198.17 (0.17) 297250 (250) $\geq$720p 25 × × 4725959 (3974) Semi-Automatic Abitrary point
TAP-Vid DAVIS [tapvid] 30 - (-) 1999 (66.6) 1080p - × × 28824 (960.8) Semi-Automatic Abitrary point
TAP-Vid RGB-Stacking [tapvid] 50 - (-) 12500 (250) 256x256 - × × 303436 (6068.7) Semi-Automatic Abitrary point
[…]
We developed a comprehensive dataset for training our unified model, termed Tracking-Any-Granularity (TAG), with annotations across three granularities: segmentation masks, bounding boxes, and key points. Our dataset contains 6,000 high-resolution videos featuring diverse scenes, objects, and challenging scenarios (e.g., occlusion, motion blur, etc.). With a three-phase data engine with model-in-the-loop annotation workflows and strict multi-stage quality checks, we ensure large-scale, high-quality, and consistent annotations. For more details, please refer to the Appendix.
5.1 Annotation Pipeline
We designed a coarse-to-fine annotation pipeline to ensure high-quality multi-granularity annotations as demonstrated in Fig. 3. Firstly, we collected videos from YouTube that meet our quality standards and exhibit diverse tracking challenges. Then comes the coarse annotation stage, where annotators mark key points and tight bounding boxes on target objects. Next, in the fine annotation stage, we leverage SAM to generate initial masks from coarse annotations, which annotators then refine. Experts perform quality checks throughout to ensure annotation consistency and accuracy, particularly for challenging scenarios like occlusions and motion blur. As for the Final Completion stage, the experts check the consistency of the three labellings.
5.2 Data Engine
As shown in Table 2, the Tracking-Any-Granularity dataset is annotated across three phases: 1) Phase ①: Manual annotation of every frame, totaling 1,000 videos. 2) Phase ②: Manual annotation of every 10 frames, totaling 2,000 videos. 3) Phase ③: Manual annotation of every 20 frames, totaling 3,000 videos. In Phases ② and ③, we integrated SAM 2++, which is trained on public datasets and previous phase annotations, to automatically annotate frames between manual-annotated frames. In detail, we divided videos into clips where both first and last frames were manually annotated, then used the annotation of the first frame in each clip as input to infer intermediate frames. To improve annotation quality, we implemented two optional enhancements: (1) performing backward tracking and fusing results with forward tracking, and (2) using the first video frame (guaranteed to contain the target) as an additional starting point when targets might be absent in keyframes. We evaluated these enhancements on validation data in Phase ① to select optimal strategies for each tracking task.
5.3 Tracking-Any-Granularity Dataset
Compared with existing datasets in video tracking tasks, our Tracking-Any-Granularity dataset stands out as the only one providing annotations at all three granularities simultaneously. We compare our dataset with numerous public datasets in Table 1, showing our dataset contains significantly more videos and annotations than they do, creating a substantial resource for multi-granularity tracking research. Fig. 4 shows examples from our dataset, annotated at all three granularities and exhibiting diverse challenges.
Table 2: Evolution of data engine phases, showing the interval and number of manual annotations.
TAG Videos Interval Points Boxes/Masks Total Frames Total Len (s)
Phase ① 1,000 1 523,137 348,715 354,625 12,540.9
Phase ② 2,000 10 87,708 75,923 787,643 28,164.5
Phase ③ 3,000 20 60,809 53,917 1,058,623 39,617.1
Total 6,000 - 671,654 478,555 2,200,891 80,322.4
Table 3: State-of-the-art comparison on Video Object Segmentation Task.
[…]
Methods Hall Hcom Hunc $\mathcal{J}\&\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}\&\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ $\mathcal{J}$ $\mathcal{J}_{tr}$ Overall $\mathcal{J}_{seen}$ $\mathcal{F}_{seen}$ $\mathcal{J}_{unseen}$ $\mathcal{F}_{unseen}$
STCN [stcn] 45.8 45.6 45.9 62.0 58.5 65.4 50.1 46.1 54.1 70.4 65.9 75.0 76.2 72.2 80.2 32.7 22.7 82.8 81.3 85.6 78.3 86.0
AOT-SwinB [aot] 54.9 55.3 54.8 73.8 70.1 77.4 60.2 56.2 64.1 78.1 73.1 83.2 80.9 76.4 85.4 39.3 29.4 85.3 84.6 89.5 79.3 87.7
DeAOT-SwinB [deaot] 58.5 58.0 58.6 72.1 68.5 75.8 61.7 57.6 65.9 79.6 74.8 84.4 81.6 77.3 85.9 43.0 29.4 86.4 85.4 90.3 80.6 89.3
XMem [xmem] 52.3 51.9 52.3 64.7 61.8 67.5 56.3 52.1 60.5 74.4 70.1 78.6 75.7 71.8 79.6 37.9 25.0 85.5 84.3 88.6 80.5 88.7
DEVA [deva] 56.1 55.4 56.2 72.2 68.6 75.9 66.4 62.2 70.6 77.9 73.1 82.6 82.1 78.0 86.1 40.9 27.3 86.3 85.3 89.8 80.6 89.2
Cutie-base+ [cutie] 55.8 55.6 55.8 71.4 68.6 74.3 66.2 62.3 70.1 79.0 75.0 83.0 83.8 80.0 87.7 44.7 32.7 86.9 86.2 90.7 81.6 89.2
[…]
JointFormer [jointformer] - - - 71.7 68.8 74.7 69.7 65.8 73.6 76.6 72.8 80.5 79.1 75.5 82.7 - - 87.0 86.1 90.6 82.0 89.5
[…]
Table 4: State-of-the-art comparison on Single Object Tracking Task.
Methods GOT10ktest $_{SOT}^{test} TAG$$_{SOT}^{val}$ TrackingNet TNL2K VastTrack
AO $SR_{0.5}$ $SR_{0.75}$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$ AUC $_{Norm} PP$
OSTrack [ye2022joint_OStrack] 74.8 84.4 72.7 69.7 78.8 69.9 68.3 77.1 66.2 83.9 88.5 83.2 57.6 74.4 58.9 33.7 40.8 31.4
SimTrack [pauwels2015simtrack] 71.1 80.5 68.1 64.1 72.4 60.5 65.8 73.7 63.7 82.3 - 86.5 54.4 70.2 53.7 34.5 40.5 30.4
[…]
DropTrack [wu2023dropmae_droptrack] 76.8 86.9 74.4 71.1 80.5 72.1 70.8 80.4 69.4 83.8 88.5 83.1 58.5 75.7 60.3 37.5 45.9 36.4
GRM [Gao_2023_CVPR_GRM] 73.1 82.3 71.4 69.1 77.4 69.1 68.5 77.3 66.5 84.0 88.7 83.3 56.8 73.2 57.7 34.6 42.2 32.3
SeqTrack [chen2023seqtrack] 77.0 85.8 76.1 69.8 79.4 71.5 68.5 78.2 67.8 83.9 88.8 83.6 57.8 75.3 60.8 35.8 44.8 35.3
ARTrack [wei2023autoregressiveARtrack] 76.8 85.8 75.7 71.1 78.7 70.9 69.9 76.9 67.1 84.2 88.7 83.5 57.9 73.9 59.6 35.7 42.1 32.4
ARTrack-V2 [bai2024artrackv2] 76.3 85.5 74.3 71.8 79.5 71.9 70.2 78.0 68.3 84.9 89.3 84.5 57.1 73.4 58.7 37.0 44.5 34.8
ROMTrack [cai2023robust_romtrack] 75.6 85.4 73.7 71.3 80.8 72.8 69.2 77.8 68.5 84.1 89.0 83.7 58.2 75.3 59.8 37.1 45.5 36.2
HIPTrack [cai2024hiptrack] 78.2 88.5 76.6 71.4 81.0 72.5 72.0 82.2 71.0 84.5 89.1 83.8 59.8 77.1 62.1 38.6 46.3 36.8
LoRAT [lin2024tracking_lorat] 75.1 84.8 74.4 70.5 79.7 68.7 72.7 82.2 74.4 83.5 87.9 82.1 58.8 76.2 61.4 38.7 41.1 37.8
[…]
Scene and Attribute. To enable a more comprehensive analysis of tracking approaches, it is critically important to identify video scenes and attributes of our dataset. Fig. 5 demonstrates that our dataset encompasses a diverse range of sources, highlighting its robust diversity and enabling it to serve as a powerful benchmark for evaluating tracking performance across various environments. Furthermore, we label each sequence with 18 attributes that represent various video challenges. It is worth noting that these attributes are not mutually exclusive, and a single video may contain multiple challenges. Fig. 5(a) and (b) illustrate the distribution of challenges in each video and their mutual dependencies. Motion Blur, Deformation, and Partial Occlusion are the most common challenges in our dataset, demonstrating its high level of difficulty. We further explore the likelihood of videos being linked to multiple attributes, and Fig. 5(c) indicates that most videos possess more than one attribute.
Dataset Splits. We selected 150 validation videos and 150 test videos from the 1,000 fully annotated videos in Phase ① with stratified sampling based on both category and source, which ensures a balanced distribution.
6 Experiments
6.1 Comparison to state-of-the-art on three tasks
Video Object Segmentation. The comparisons between our model and previous semi-supervised VOS methods are demonstrated in Table 3, including YoutubeVOS-19 [youtube-vos], MOSE [MOSE], LVOS-v2 [LVOS_V2], BURST [burst], VOST [vost], VISOR [visor], and our TAG dataset. We use the standard metric $\mathcal{J\&F}$ [vos-benchmark] that averages Jaccard index and contour accuracy in most benchmarks, but adopt Higher Order Tracking Accuracy (HOTA) [hota] in the BURST benchmark. Results show that our model outperforms individual video object segmentation models.
Single Object Tracking. We compare the performance of our proposed model on three benchmarks in Table 4, including TrackingNet [TrackingNet], GOT-10k [got10k], TNL2K [wang2021tnl2k], VastTrack [vasttrack] and our TAG dataset, and all compared models are trained on four datasets. We choose the Average Overlap (AO) for the GOT-10k benchmark, and Area Under the Curve (AUC) for the other benchmarks. Experiments demonstrate that our model consistently outperforms existing methods across nearly all benchmarks.
Table 5: State-of-the-art comparison on Point Tracking Task. model and model means offline and modified online trackers.
Method Key Point Arbitrary Point
BADJA PerceptionTestval $_{PT}^{test} TAG$$_{PT}^{val}$ Tapviddavis Tapvidrgb
pips [harley2022pips] 64.2 41.5 19.0 19.8 40.9 28.5
pips [harley2022pips] 56.9 27.2 12.3 12.2 26.3 20.8
pips++ [zheng2023pointodyssey] 56.6 62.6 20.9 23.1 59.8 58.5
CoTracker [cotrackerv1] 65.3 59.8 23.3 22.3 60.9 63.1
CoTracker [cotrackerv1] 55.1 48.9 18.8 18.1 50.8 46.1
CoTracker3 [cotrackerv3] 72.7 71.3 29.6 29.1 65.6 70.6
CoTracker3 [cotrackerv3] 66.3 66.3 25.8 24.9 59.2 63.6
TAPTR [TAPTR] 69.1 59.4 23.7 23.8 61.2 58.0
TAPTR [TAPTR] 63.0 48.9 20.4 19.0 52.3 39.4
TAPIR [doersch2023tapir] 64.2 59.6 21.3 24.6 56.8 50.4
LocoTrack [LocoTrack] 68.7 67.1 25.2 30.2 62.7 69.2
Track-On [Track-On] 69.7 69.7 24.8 25.8 64.5 64.5
[…]
Figure 6: Visualization of memory design at different components and granularities. In the visualization of each component, the left side represents our task-adaptive memory mechanism and the right side represents full parameter sharing.
Online Point Tracking. We compare our method to prior works in Table 5 on four benchmarks, including BADJA [badja] and Perception Test [patraucean2023perception] for key point tracking, TAP-Vid [tapvid] for arbitrary point tracking, and our TAG dataset in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries. We report the Percentage of Correct Keypoint-Transfer (PCK-T) for the BADJA benchmark, and Average Jaccard (AJ) for the remaining benchmarks. However, most of the current methods are offline trackers, which process long-temporal window frames or even the entire video to be able to see the future frame, and do not match the online setup required by real applications. For a fair comparison, we modified their input so that there is no future information inside the window, denoted as model, to enable inference online. Experimental results show a substantial decrease in model performance when the input data is switched from offline to online. The comparative analysis reveals the effectiveness of our approach on keypoint tracking benchmarks, which surpasses competing models. Furthermore, although our model is trained on keypoint datasets, it demonstrates generalization capability on arbitrary point tracking datasets.
Table 6: Analysis of Mixed training strategy and Data Engine. ‘Shared’ and ‘Decoupled’ denote that the module shares parameters or decouples parameters between different tasks, respectively
[…]
$\times$ Shared Shared 69.3 84.7 28.1 26.1 64.1 30.3
$\times$ Shared Decoupled 73.6 86.4 64.5 57.3 63.0 28.7
+① Freezed Decoupled 74.2 87.1 67.8 59.2 68.8 29.9
+① Decoupled Decoupled 75.4 86.8 66.9 58.5 72.1 32.7
+① Shared Decoupled 76.4 87.1 68.9 58.2 71.9 33.1
[…]
6.2 Exploration Studies
Study on Mixed training strategy. To verify the effectiveness of task-adaptive memory during multi-task joint training, we compare the results of single-task training with different parameter settings during multi-task mixing training. As shown in Table 6, when a single set of parameters is naively shared for multi-task joint training, the differences between tasks lead to a performance drop across all tasks. This indicates that the encoding and retrieval components of the memory module need to be decoupled for different tasks. In addition, when the image encoder is either frozen or similarly decoupled, the performance is inferior to a shared encoder. This suggests that the image encoder benefits from exposure to more data and is not adversely affected by task differences.
Study on data engine. To validate the effectiveness of our data engine, we evaluated the performance when trained on different phases of our TAG dataset, as shown in Table 6. The results demonstrate that our proposed dataset enhances the performance on other datasets, indicating high diversity and generalizability. After training with more data from phases ② and ③, the performance is further improved across all three tasks, demonstrating the effectiveness of the supplementary data provided by our data engine.
6.3 Visualization
To further illustrate the varying requirements for memory representation of targets at different granularities, we visualize the memory-related outputs for the three tasks under both task-adaptive memory mechanism (left) and full parameter sharing (right), as shown in Fig. 6. Firstly, we observe that even under different training settings, the memory features for the same task remain highly similar, indicating that different granularities have distinct memory requirements. Secondly, under the memory-related decoupled training setting with our task-adaptive memory mechanism, both memory attention and mask output align more closely with the task outputs compared to full parameter sharing, highlighting the necessity of the decoupled design. Finally, we find that under the full parameter sharing setting, the mask output for point tracking does not exhibit a Gaussian pattern, leading to incorrect predictions, while the training method with a decoupled design does not exhibit this error. This demonstrates that the decoupled design effectively preserves the specific needs of different tasks.
7 Conclusion
We present SAM 2++, a foundational model for tracking targets at any granularity, built upon three key contributions: 1) Unifying task processing through task-specific prompts for inputs and a Unified Decoder for outputs; 2) Unifying task states across different granularities via a task-adaptive memory mechanism; 3) Introducing the Tracking-Any-Granularity dataset for training and benchmarking video tracking at multiple granularities. We hope that SAM 2++ can serve as a strong baseline for general tracking and provide a powerful impetus for future research.
[…]
Supplementary Material
In the appendix, we present a more detailed discussion of the topics covered in the main text, with the specifics of each section described as follows:
  • Section 8: Model Details
  • Section 9: Data Details
  • Section 10: Additional Experiments
  • Section 11: Limitations, Impacts, and Future
8 Model Details
8.1 Model Architecture
Task-Specific Prompt. In order to unify the different inputs for each task and not modify the structure of the original Prompt Encoder, we provide task-specific prompt for each task, which provides an accurate and efficient representation of the target state of each task. The design of the task-specific prompt for each task is as follows:
  • Mask tracking: {0, 1} mask to accurately describe the shape and boundaries of the target;
  • Box tracking: bounding boxes in the top-left and bottom-right corners;
  • Point tracking: Besides the exact point coordinates, we provide a (0, 1) Gaussian mask generated from the points to better represent the target in memory and align with the mask outputs from the Decoder.
Unified Decoder. We made minor modifications to the Mask Decoder to obtain the desired outputs for each task. Specifically: (1) we added a Corner Head for the box tracking task to directly output the bounding box, thereby avoiding the low precision and lack of gradients associated with the outer box operation; and (2) we applied an argmax operation to the masked output (or a soft-argmax operation during training to ensure gradient flow) to obtain the point coordinates, which are aligned with the Gaussian mask of the prompts.
Task-adaptive Memory. Based on the analysis described in the main text, we decouple the memory components for each task. Specifically, for Memory Attention in the Transformer architecture, we configure an independent LoRA for each task. For the Memory Encoder in the convolutional architecture, we set up a separate copy for each task, resulting in only a minimal increase in parameters.
8.2 Training and Inference Details
Table 7: Hyperparameters and details of SAM 2++ training in three tasks.
Settings mask box point
dataset
DAVIS 2017,
MOSE,
YoutubeVOS 2019 ,
Tracking-Any-Granularity
LaSOT,
COCO,
GOT-10K,
TrackingNet,
Tracking-Any-Granularity
TAP-Vid Kinetics,
PerceptionTest,
PointOdyssey ,
Tracking-Any-Granularity
sample prob 0.1 0.4 0.5
batch size 16
drop path 0.2 (B+)
epochs 150
resolution 1024
precision bfloat16
optimizer AdamW
optimizer momentum $\beta_{1}$, $\beta_{2}$ =0.9, 0.999
gradient clipping type: L2 , max: 0.1
weight decay 0.1
learning rate (lr) backbone: 5.0e-6, other: 3.0e-06
lr schedule cosine
layer-wise decay 0.9 (B+)
video augmentation
hflip,
affine (deg: 25, shear: 20) in mask, w/o affine in box and point,
resize to 1024 (square),
colorjitter (b: 0.1, c: 0.03, s: 0.03, h: null),
grayscale (0.05),
per frame colorjitter (b: 0.1, c: 0.05, s: 0.05, h: null)
mask losses focal (20), dice (1) focal (0.5), dice (0.1) focal (20), dice (1)
task-specific losses - ciou loss (1), box IoU L1 loss (1) point distance L1 loss (20)
Iou loss l1 loss (1)
occlusion loss cross-entropy (1)
input prompt
mask (0.5),
noisy box (0.25),
sampled points (0.25)
box (0.5),
noisy box (0.25),
sampled points (0.25)
point coordinate & Gaussian mask (1)
# max. object per frame 3 1 3
# training frames 8 8 8
[…]
# corrective points 7 7 -
# num_maskmem 7 7 7
Task-specific Modification
Decoupled & LoRA Memory Attention, Decoupled Memory Encoder,
Shared & LoRA Image Encoder
Other Setting - Corner Head
Gaussian mask settings
[…]
Training. The SAM 2++ training is conducted on 16 H800 GPUs. We expand SAM 2++ to three tasks: semi-supervised video object segmentation (mask), single object tracking (box), and online point tracking (point), by processing input, prompt, memory, and output into a unified format used by SAM 2. Our training process is based on the mask tracking task in SAM 2 and we make minimal modifications to it while adding task-specific requirements from the other tasks. Table 7 describes the training settings for three tasks in detail, and other settings not mentioned follow SAM 2.
Training is performed jointly on data of the three tasks. In addition to our Tracking-Any-Granularity dataset, we used DAVIS-17 [davis17], YoutubeVOS-19 [youtube-vos] and MOSE [MOSE] for the mask task, LaSOT [lasot], GOT-10k [got10k], TrackingNet [TrackingNet] and COCO [coco] for the box task, and TAP-Vid Kinetics [tapvid], PointOdyssey [zheng2023pointodyssey], and PerceptionTest [patraucean2023perception] for the point task. To enable the model to be simultaneously capable of all three tasks and to optimize training efficiency, we adopt the strategy of alternating between the three tasks. Specifically, we implement parallelisation by sampling a whole batch at each step of training, which is entirely derived from the data of a particular task. The sampling probability is set to 1:4:5 to balance the performance of the three tasks.
We sample 8 frames from each video as a training sequence, randomly choose up to 3 targets (or 1 target box in box tracking) from the objects of this video, and ensure that these sampled targets are visible in the first frame of the sequence. We randomly select up to 2 frames from the sequence, including the first frame, as conditional frames to give these frames initial prompts. Since we prefer to maintain the interactive capabilities of SAM 2, we keep the interactive prompts in mask tracking and box tracking during training. Specifically, we start by deciding whether the conditional frames accept normal or interactive input in this training step with 50% probability: for normal input, we use ground-truth as initial prompts; for interactive input, we use a noisy bounding box or a positive click from the ground-truth with 50%-50% probability. Alternatively, suppose we use the normal input prompts in conditional frames. In that case, we directly convert them into memory instead of prediction and do not supervise their predictions for this input. However, the point tracking task requires precise inputs, so we can only provide GT points in the first frame instead of various formats of prompts like the other two tasks. As for the multi-prediction scenario, when a frame receives no prompt, or at most 1 point (the box prompt can be seen as 2 points), the model will output 3 task predictions and their iou predictions for that frame.
In addition, if interactive input is used as initial prompt, we select up to 2 frames as corrective frames to add corrective clicks on them: after predicting the selected frame, we sample a positive point from the false positive region between the prediction and ground truth or a negative point from the false negative region as a corrective point, and use it as additional prompt to get a new prediction along with all previous cumulative prompt from that frame. This operation is repeated until 7 corrective points have been added. In addition, if the box tracking task uses the box format to compute the regional differences between the prediction and the GT, there is an overwhelming problem that the sampled corrective clicks may fall at the boundaries of the box instead of inside the target, which is contrary to the actual interaction. Therefore, we choose to compute the difference in mask format, and use SAM 2 and sam-hq [sam_hq] with box annotations to obtain the pseudo-GT mask on SOT datasets because of its good segmentation ability.
Losses and optimization. Following the mask tracking task in SAM 2, we adopt the linear combination of focal loss $\mathcal{L}_{mask}^{focal}$ and dice loss $\mathcal{L}_{mask}^{dice}$ for the mask prediction, L1 loss for the IoU prediction $\mathcal{L}_{L1}^{IoU}$, and cross-entropy loss for object occlusion prediction $\mathcal{L}_{CE}^{obj}$. During the box tracking task, we adopt the corner head to predict the bounding boxes and add additional ciou loss [zheng2020diouciou] and L1 loss to supervise the box prediction. As for the point tracking task, we select the highest probability position from the mask prediction as point prediction, and use soft argmax [softargmax] during training for making the process derivable instead of the undifferentiable argmax function. Beyond the loss on mask in the form of Gaussian map, we add an L1 loss between the prediction and ground-truth point to directly optimize the distance and accuracy of the points. For multi-prediction scenario, we only supervise the task predictions (masks, boxes, and points) with the lowest loss, which is a combination of $\mathcal{L}^{mask}$, $\mathcal{L}^{box}$ and $\mathcal{L}^{point}$, but supervise the IoU predictions of all task predictions to learn to synchronise the quality of predictions. Furthermore, if the target is missing in some frames due to disappearance or cropping, we do not supervise the task predictions or iou predictions on them in all three tasks, but always supervise the occlusion prediction from an MLP head, no matter if the ground-truth exists or not. In summary, the supervision losses for the three tasks can be written as:
$\mathcal{L}_{Mask}$ $=\mathcal{L}_{mask}+\mathcal{L}_{IoU}+\mathcal{L}_{obj}$ (7)
$=\left[\lambda_{mask}^{focal}\mathcal{L}_{mask}^{focal}+\lambda_{mask}^{dice}\mathcal{L}_{mask}^{dice}\right]\times\mathbbm{1}_{obj}$
$+\lambda_{IoU}^{L1}\mathcal{L}_{IoU}^{L1}\times\mathbbm{1}_{obj}+\lambda_{obj}^{CE}\mathcal{L}_{obj}^{CE},$
$\mathcal{L}_{Box}$ $=\mathcal{L}_{Mask}+\mathcal{L}_{box}$
$=\mathcal{L}_{Mask}+\left[\lambda_{box}^{ciou}\mathcal{L}_{box}^{ciou}+\lambda_{box}^{L1}\mathcal{L}_{box}^{L1}\right]\times\mathbbm{1}_{obj},$
$\mathcal{L}_{Point}$ $=\mathcal{L}_{Mask}+\mathcal{L}_{point}$
$=\mathcal{L}_{Mask}+\lambda_{point}^{L1}\mathcal{L}_{point}^{L1}({GT}_{point},{O}_{point})\times\mathbbm{1}_{obj},$
where $\mathbbm{1}_{obj}$ denotes we supervise task and IoU prediction only if the object exists, and $\lambda$ represents the weights of different losses. The specific hyperparameters for the training are shown in Table 7.
Inference. We conduct all benchmarking experiments on a single A100 GPU using PyTorch 2.5.1 and CUDA 12.1, under automatic mixed precision with bfloat16. We inference all three tasks following the fully online inference setting, i.e., all operations in the current frame can not see the future and only the ground-truth in the first frame is given as a prompt for each target object at the beginning of the sequence without any correction input in the subsequent frames. For mask tracking task (VOS), we first give each object the ground-truth mask in the first frame and make mask predictions for each object independently and in parallel. In the multi-object scenario, we merge the per-object logits into a single mask by simply fusing the mask logits based on their values. For the box tracking task (SOT), the bounding box prediction of the object can be obtained directly from the corner head.In case of the point tracking task, we replace the prompt with the ground-truth point coordinates and an additional generated mask in Gaussian form, and use the argmax operation to obtain the point coordinates from the mask prediction. Note that our model is a neat tracker where inference is performed on the complete current frame without any post-processing strategies. For example, the centre crop operation, a widely used operation in the SOT task, is able to pre-crop the current frame according to the location in the previous frame, avoiding some incorrect tracking.
8.3 Efficiency Analysis
Table 8: Hyperparameters and details of SAM 2++ training in three tasks.
Modules GFlops Main Param Lora Param
Image Encoder 264.4 69.1 22.25
Memory Encoder 5.0 1.4 $\times$ 3 -
Mask Decoder 53.4 9.9 -
Memory Attention 27.4 5.9 4.3 $\times$ 3
Total 350.2 89.1 35.1
To present a comprehensive view of the model’s computational complexity and parameter overhead, we provide a detailed breakdown of the computational cost for each module in Table 8, including GFLOPs, the number of parameters, and the LoRA parameters introduced during training. We would like to kindly note that the LoRA parameters exist only during training and are merged into the main model weights at inference time. Therefore, they do not introduce additional parameters or computational overhead during inference. In addition, although the model contains multiple Memory Encoders designed for different granularities, only the branch corresponding to the current granularity is activated at inference, while the others remain inactive (excluded from computation), ensuring that no extra inference cost is incurred.
9 Data Details
The key features of this dataset are as follows: (1) High Resolution: The dataset consists of high-resolution videos, ensuring that fine details are preserved and enabling more accurate analysis. (2) Diversity: It encompasses a wide variety of scenes, sources, and tracked object categories, providing a rich and representative sample of real-world scenarios. (3) Complex and Challenging Cases: The dataset includes numerous complex situations, such as occlusion, motion blur, and other challenging visual conditions, which test the robustness and generalization ability of tracking algorithms. (4) Comprehensive Annotations: the dataset contains annotations at multiple granularities, including segmentation masks, bounding boxes, and key points.
9.1 Data Requirements
Videos Requirements. The selected videos must satisfy the following criteria:
  • No camera cuts or scene transitions are present throughout the video.
  • Visuals are clear, and the boundaries of the target can be accurately identified.
  • The duration is between 10 and 40 seconds (excluding static images).
  • Each video must contain at least one target object that meets the outlined below.
Target Object Requirements. Each video must include at least one object designated as the tracking target, which must fulfill the following basic criteria:
  • The target has clearly distinguishable boundaries from other objects in the scene.
  • Eligible targets include the full body or parts of a human (e.g., face, facial features, limbs, hands, feet, etc.) or an animal (full body or parts).
  • The target must appear in the first frame of the video and be clearly identifiable.
  • At least one key point on the target must be visible and locatable for most of the video, allowing brief occlusions or exits.
  • The target should be in motion (either actively or passively) for most of the video.
  • To ensure the dataset emphasizes challenging tracking scenarios, the target must also meet at least one of the following additional difficulty criteria:
    • Rapid movement of the target itself or due to camera motion.
    • High similarity to other objects.
    • Occlusion or brief disappearance and reappearance.
    • Deformation (e.g., shape or structure changes) or notable changes in size, orientation, or viewpoint (e.g., approaching or turning).
    • The target is small relative to the frame, but not excessively tiny.
Target Point Requirements. We further pick at least one point on the chosen target object. These points need to meet the following conditions:
  • The point could be the center point, the corner point, or semantically meaningful points such as human eyes, hands, or head.
  • The key point must be present in the first frame.
  • If the key point is occluded or disappeared, it should be labeled as ”occluded.”
  • For spherical objects, the key point should be placed near the center.
Many current point tracking datasets use arbitrary points as annotation targets. However, we chose to focus on keypoints as the target for both data annotation and model optimization for the following reasons, primarily based on two considerations: 1) Practical Application Perspective: Downstream tasks like 3D reconstruction and SLAM, require tracking key points in most cases. Key points offer stronger distinguishing and descriptive capabilities, and typically only a small number of high-quality key points are sufficient for other tasks, eliminating the need to track any-point. 2) Annotation Cost Efficiency: Annotating any-point incurs prohibitively high costs. Unlike RoboTAP (which relies on optical-flow-based trajectory interpolation and is limited to lab scenes) or Kubric/RGB-Stacking (which generates point annotations via rendering, lacking real-world diversity), our dataset sources videos from indoor, outdoor, and wild environments from real-world. To ensure annotation accuracy, the target points were selected by annotators and manually labeled frame-by-frame. Due to the unbearable time and human resources required for any-point annotation, keypoint annotation is a better choice to balance dataset utility and feasibility. Similarly, real-scene datasets like DAVIS and Kinetics annotate most salient objects.
Table 9: Performance Comparison of Automatic Visible Annotation.
Type + 0th Backward Visible Acc.val Precisionval Recallval F1val Acc.test Precisiontest Recalltest F1test
#01 AND 98.47 99.52 98.86 99.12 98.89 99.54 99.27 99.38
#02 OR 98.84 98.85 99.90 99.35 99.01 99.06 99.89 99.45
#03 SOT 98.19 98.96 99.13 98.96 98.68 99.10 99.50 99.26
#04 VOS 99.12 99.42 99.63 99.51 99.22 99.49 99.66 99.56
#05 AND 98.72 99.65 98.98 99.27 98.94 99.70 99.15 99.40
#06 OR 98.71 98.65 99.98 99.27 98.91 98.96 99.90 99.40
#07 SOT 98.36 98.70 99.58 99.07 98.80 98.96 99.78 99.34
#08 VOS 99.08 99.60 99.38 99.48 99.05 99.70 99.27 99.46
#09 AND 98.49 99.52 98.89 99.13 98.96 99.56 99.32 99.42
#10 OR 98.84 98.83 99.92 99.34 99.01 99.04 99.91 99.45
#11 SOT 98.21 98.94 99.17 98.97 98.68 99.08 99.52 99.26
#12 VOS 99.12 99.41 99.63 99.51 99.30 99.51 99.72 99.61
#13 AND 98.79 99.69 99.02 99.31 99.03 99.70 99.23 99.45
#14 OR 98.73 98.67 99.98 99.28 98.91 98.94 99.92 99.40
#15 SOT 98.37 98.72 99.57 99.08 98.78 98.94 99.78 99.32
[…]
Table 10: Performance Comparison of Automatic Annotation in Different Annotation Methods.
(a) mask automatic annotation
Type + 0th Back. $\mathcal{J}\&\mathcal{F}^{val}$ $\mathcal{J}^{val}$ $\mathcal{F}^{val}$ $\mathcal{J}\&\mathcal{F}^{test}$ $\mathcal{J}^{test}$ $\mathcal{F}^{test}$
#1 94.4 91.7 97.1 94.6 91.9 97.3
#2 95.0 92.4 97.6 95.0 92.4 97.6
#3 94.5 91.8 97.2 94.7 92.0 97.4
[…]
(b) box automatic annotation
Type + 0th Back. AUCval $P_{Norm}^{val}$ $P^{val}$ AUCtest $P_{Norm}^{test}$ $P^{test}$
[…]
#2 83.7 93.9 91.3 84.5 94.7 91.0
#3 83.9 93.8 91.3 84.7 94.5 91.0
#4 83.6 93.6 91.1 84.4 94.4 90.9
(c) point automatic annotation
+ 0th Back. Vis. OAval AJval OAtest AJtest
#1 - 89.0 61.7 88.7 61.3
[…]
#3 OR 89.3 58.1 89.6 57.3
#4 AND 88.2 62.1 87.7 61.1
#5 OR 89.4 58.8 89.6 57.7
#6 AND 88.4 62.6 87.8 61.4
Figure 7: Example videos from the Tracking-Any-Granularity dataset with annotation at various granularities. Each annotation has a unique color. Better viewing with zoom and color.
9.2 Annotation Pipeline
We designed a coarse-to-fine annotation pipeline to ensure high-quality multi-granularity annotations, which consists of the following four steps.
1) Video Selection. We downloaded a large number of videos from YouTube and instructed the annotators to select videos and objects that meet the above requirements.
2) Coarse Annotation. Annotators mark key points and tight bounding boxes on target objects.
3) Fine Annotation. To reduce annotator workload and improve efficiency, we use SAM [kirillov2023seganysam1] to generate rough masks based on the coarse annotations (points and boxes). Then, annotators refine these masks with the following requirements:
  • Only annotate the visible parts of the present object.
  • In cases of motion blur, infer the approximate position based on the previous frame to maintain temporal consistency. Masks in adjacent frames should not differ drastically.
  • Ignore transparent or semi-transparent watermarks and subtitles when creating masks; masks can directly cover these elements.
  • Exclude opaque overlays (such as logos or captions) from the mask.
  • For containers holding other objects, do not include the contained objects in the mask.
  • The mask should tightly fit the object, neither exceeding nor falling short of its boundaries.
  • Ensure that mask edges are smooth and avoid excessive roughness.
  • Fill in small internal holes, but preserve natural gaps (such as hollowed-out structures) or occlusions caused by other objects.
  • If the initial SAM-generated mask is of very poor quality, annotators may clear it entirely and use color tolerance-based selection to manually annotate the object from scratch.
4) Final Completion. Experts perform a final review to thoroughly assess the accuracy and consistency of all three types of annotations, ensuring that the labeling meets the required standards and that any discrepancies are identified and corrected.
9.3 Data engine
To increase the size of the dataset while reducing the workload, we adopted a selective annotation strategy in the second and third phases. Instead of manually labeling every video frame, annotators labeled only a subset of frames at varying intervals. After training the model on both public datasets and the fully labeled data from earlier phases, we leveraged the model to automatically annotate the remaining frames. Specifically, each video was divided into multiple clips, with annotators manually labeling the first and last frames of each clip. The annotation of the first frame in each clip served as the initial target state, enabling the model to infer the target state in the intermediate frames.
To further enhance annotation quality, we introduced two optional refinement methods: (1) performing backward tracking and fusing the results with those from forward tracking, and (2) since the target may be absent in some annotated frames, using the first frame of the entire video, which is guaranteed to contain the target, as an additional reference state alongside the first frame of each clip. We evaluated these enhancement methods on the Phase 1 validation and test set to determine the inference setting for each tracking task, as shown in Table 10. Specifically, (a), (b), and (c) represent the evaluation outcomes for the VOS, SOT, and PT tasks under various settings, respectively, while Table 9 shows the evaluation results for object existence prediction. By comparing the results across different settings, we select the configuration highlighted in the gray row as the inference setting for each task.
Table 11: Statistical analysis of video data from our dataset.
Average Medium Minimum Maximum
Frame per Video 366.8 295 80 3,317
Length(s) per Video 13.39 10.9 5.3 110.9
Video FPS - 24 10 60
Table 12: Attribute analysis of video data from our dataset.
Attribute. Definition Num.
BC. Background Clutter. The appearances of background and target object are similar. 1284
CS. Camera-Shake. Footage displays non-negligible vibrations. 429
DB. Dynamic Background. Large movement of background areas or other objects. 1523
DEF. Deformation. Target appearance deform complexly. 3580
EA. Edge Ambiguity. Unreliable edge detection, such as thorny sea urchins and rolling waves. 339
FO. Full Occlusion. Object becomes fully occluded, accompanied by Partially Occlusion in most cases. 253
FOV. Fully Out of View. The object is fully clipped by the image boundaries. 203
IV. Illumination Variation. when illumination in object region heavily varies. 232
MB. Motion Blur. Boundaries of target object is blurred because of camera or object fast motion. 4332
OP. Object Part. The object is a part of the whole. 761
PO. Partially Occlusion. The object becomes partially occluded. 2844
POV. Partially out of view. The object is partially clipped by the image boundaries. 1498
ROT. Rotation. The object rotates. 2279
$R$. Reflected and Refraction. The object undergoes reflection or refraction. 267
SA. Similar Appearance. There are multiple different objects that are similar to the target object. 1546
SC. Shape Complexity. Boundaries of target object is complex. 123
TO. Transparent Object. The object is transparent 87
VC. Viewpoint Change. The camera viewpoint changes. 1304
9.4 Tracking-Any-Granularity Dataset
Our dataset comprises 6,000 videos, each annotated with three types of labels: masks, boxes, and points. Fig. 7 shows some videos with various annotations.
Statistics and Attribute. The resolution of the majority of the videos is $1280\times 720$, with 398 exceptions. The duration of the videos ranges from $5.3$ seconds to $110.9$ seconds, and the frame count varies from $80$ frames to $3,317$ frames. In total, the dataset comprises $2.2$ million frames, amounting to a cumulative duration of $1,338.7$ minutes. More detailed statistics are shown in Table 11. We label each sequence with 18 attributes that represent various video challenges, as shown in Table 12.
10 Additional Experiments
10.1 Performance Comparison
Evaluation metrics. In video object segmentation task, we use standard metrics [vos-benchmark] in most benchmarks: Jaccard index $\mathcal{J}$, contour accuracy $\mathcal{F}$, and their average $\mathcal{J\&F}$. In the YouTubeVOS benchmark, $\mathcal{J}$ and $\mathcal{F}$ are computed for ”seen” and ”unseen” categories separately. $\mathcal{G}$ is the averaged $\mathcal{J\&F}$ for both seen and unseen classes. In LVOS benchmark, the first densely annotated long-term VOS dataset with high-quality annotations, it introduces the standard deviation $\mathcal{V}$ of the average score of $\mathcal{J}$ and $\mathcal{F}$ to assess the temporal stability of VOS models. In VOST benchmark, which focuses on segmenting objects as they undergo complex transformations, it additionally reports $\mathcal{J}_{tr}$ for the last 25% of the frames in a sequence to show the robustness after the transformation has been mostly completed. The BURST benchmark is evaluated with Higher Order Tracking Accuracy (HOTA) [hota] as a good balance between measuring frame-level detection and temporal association accuracy. In single object tracking task, we evaluate performance with Area Under the Curve (AUC), normalized precision (${P}_{Norm}$) and precision (P{P}) to measure the average accuracy of center, size and scale between the prediction and labeled groundtruth bounding boxes of all the frames for most benchmarks. For the GOT-10k benchmark, we choose the average overlap (AO) and success rate (SR) as indicators. The former AO denotes the average of overlaps between all groundtruth and estimated bounding boxes, while the SR measures the percentage of successfully tracked frames where the overlaps exceed a threshold (e.g., 0.5). In point tracking task, we report Occlusion Accuracy (OA) and Average Jaccard (AJ) for TAP-Vid [tapvid], Perception Test [patraucean2023perception], and our dataset. As for the BADJA benchmark, we adopt the Percentage of Correct Keypoint-Transfer (PCK-T). We measure these benchmarks in the ‘query first’ evaluation, which means points appearing in the first frame are used as queries.
Figure 8: Examples from SAM 2++ results on video benchmarks at various granularities.
Figure 9: Comparison between our model and various SOTA methods on video tracking benchmarks at three granularities. Better viewing with zoom and color.
Table 13: Comparison with Unified Models.
[…]
$\mathcal{J}\&\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ AUC $_{Norm} PP$
Unicorn [Unicorn] 69.2 65.2 73.2 83.0 86.4 82.2
UNINEXT-H [UNINEXT] 81.8 77.7 85.8 85.4 89.0 86.4
MITS [MITS] 84.9 82.0 87.7 83.4 88.9 84.6
OmniTracker-L [OmniTracker] 71.0 66.8 75.2 83.4 86.7 82.3
[…]
Comparison with Unified Models. To demonstrate the superior performance of our unified tracking model, we conducted a comparative evaluation against other unified models. As shown in Table 13, our model achieves significantly better results on two classical benchmark datasets, highlighting its remarkable effectiveness and robustness.
Table 14: Comparison with TAG Model.
Methods VOST LaTOT TOTB CroHD
TAG [harley2024tag] 31.3 35.3 74.4 57.1
[…]
Comparison with TAG model. We found that there exists a work, TAG [harley2024tag], that shares the same objective as ours, which is to achieve tracking of mask, box, and point with a unified model. To highlight the contribution of our work, we provide a detailed comparison between the two approaches. Firstly, TAG is an offline tracking model that processes multiple frames as a clip simultaneously, which not only differs from the current mainstream online frame-by-frame tracking pipeline but also leads to information leakage from future frames and is only applicable to pre-recorded videos. In contrast, our model ensures that the current frame only receives information from the past frames, making it suitable for video streams. Secondly, in the way of the prompt construction, TAG simply converts point coordinates into a {0, 1} mask in the point task, providing limited target information. Our method combines point coordinates with a (0, 1) Gaussian mask: the former provides precise locations, while the latter highlights the target point in mask form, maintaining consistency with output from MaskDecoder and input for MemoryEncoder, thereby enhancing expressiveness. For the box task, TAG converts the box into a square mask, which causes confusion between the target region and the background, affecting the accuracy of the target information. Third, TAG is trained only on public datasets, which limits the scale of the training dataset. In contrast, we construct a Data Engine that enables both model training and dataset annotation expansion, ultimately resulting in a large-scale dataset with three types of granularity annotations and a well-trained model. Most important of all, as an offline approach, the TAG model primarily focuses on how to jointly encode targets of varying granularity. When processing the next clip, the prompt remains in an original, unmodeled form, lacking rich target representation (e.g., mask, point, or box). Meanwhile, due to the lack of judgment of predictions, the next clip must adopt the prediction of the last frame in the previous clip as a prompt, even if it may be unreliable, which leads to error accumulation and makes it difficult to handle common challenges such as temporary target disappearance. In contrast, our method follows the online setting. The core challenges lie not only in multi-granularity prompts encoding, but also in how to transform predictions into memory representations to guide subsequent frames. Compared to original prompts, memory offers richer target features that improve the stability and accuracy. Leveraging both selective capability and memory diversity, the mechanism compensates for potential errors in individual predictions, effectively improving stability. To make this key component compatible with multi-granularity tracking, we introduced a Task-Adaptive Memory mechanism, one of the major contributions, to unify the predictions of varying granularity into memory representations. In summary, the distinction is not merely at the task level, but it directly impacts motivation and core innovations. Additionally, we present a performance comparison on the three tasks. As shown in the Table 14, our model significantly outperforms the TAG model on all three tasks, demonstrating the superior performance of our model.
Table 15: Analysis of training data and task mixtures on three tracking tasks.
[…]
$\mathcal{J}\&\mathcal{F}$ $\mathcal{J}$ $\mathcal{F}$ AUC $_{Norm} PP$ AUC $_{Norm} PP$
SAM 2 - - 73.6 77.6 69.5 82.0 92.2 81.6 65.4 72.6 69.7 $\times$ $\times$
#1 - - - 86.3 94.7 87.3 68.7 75.4 73.5 66.2 78.6
#2 74.4 70.4 78.4 85.8 94.1 86.5 68.8 75.7 73.5 63.0 71.3
[…]
Qualitative Results. We first demonstrate the multi-granularity tracking capabilities of SAM 2++ across multiple benchmarks, as illustrated in Fig. 8. We further compare our method qualitatively with various SOTA models in three tasks. As shown in Fig.9, our model outperforms other models across all three tasks. This demonstrates that our model effectively handles various target state granularities while also exhibiting strong robustness and generalization to diverse scenarios and challenges.
10.2 Model and Data Ablation
Study on model setting of point tracking task. We compare the performance of point tracking under different model settings in Table 16. First, performance declines when the Gaussian mask prompt for the point tracking task is removed, indicating that incorporating the Gaussian mask effectively assists the mask output of the Decoder, and demonstrating the effectiveness of our proposed task-specific prompt. Second, we compare two approaches for obtaining point coordinates: applying argmax to the mask output v.s. adding an MLP to predict the coordinates directly. The results show that the argmax operation yields better performance, suggesting that argmax is an effective method for point prediction, supervises the mask output as a memory source, and better represents the target state at the point granularity. In contrast, the additional MLP requires adaptation to the original model and struggles to supervise the mask output effectively.
Table 16: Analysis of the model setting on point tracking task.
Type prompt coordinates BADJA
#1 Coord. argmaxargmax(mask) 65.6
#2 Coord. & Gauss. Mask MLP 64.8
[…]
Study on task mixture and training data. We compare the performance under different training settings for three tracking tasks in Table 15. For evaluating original SAM 2 on the single object task, we take the ground-truth bounding box from the first frame as a box prompt to predict the target mask, then predict the mask frame by frame, and finally extract the outer bounding box from each mask as the final box prediction. After training on the public dataset and Phase ① of our Tracking-Any-Granularity dataset, the performance of our SAM 2++ model improves across all three tasks, demonstrating the advantages of our model design. More importantly, when we further incorporate two additional tasks during training, the model’s performance on both tasks surpasses that of training on a single task alone. This illustrates two core motivations behind our proposed model: (1) Although the granularity of the target states in the three tasks differs, they all can adopt the ”matched memory” tracking paradigm. Thus, training on various tasks enhances the matching ability, which in turn improves the performance of all tracking tasks. (2) As a generalized model supporting multiple tasks, SAM 2++ can be trained on large-scale datasets for multiple tasks, rather than being restricted to individual tasks. Finally, under the task-mixed training setting, incorporating our proposed dataset further improves the model performance on both tasks. This improvement demonstrates that the diverse and comprehensive annotations included in our dataset provide valuable supervision signals for the model, enabling it to learn more robust and generalizable representations.
11 Limitations, Impacts, and Future
As a foundational model, SAM 2++ demonstrates strong performance in video tracking tasks across all three granularities, setting a new and powerful benchmark in the field of general video tracking. As an annotation tool, SAM 2++ supports tracking multi-granularity, which greatly reduces the time and cost required to switch trackers between different application scenarios. Furthermore, its ability to automatically generate annotations at multiple granularities provides an efficient and accurate tool platform for a wide range of research fields.
However, the model still has some limitations. First, the current version does not yet support language- and audio-based references. Addressing this limitation requires integrating corresponding feature extractors into the Prompt Encoder to accommodate more types of reference states, as well as introducing relevant datasets for training. Second, in our task-specific memory, some parameters of the memory-related modules are decoupled for different tasks. Although this mechanism only adds a minimal number of parameters, these parameters are supervised by a single task and cannot benefit from multi-task learning as the majority of shared parameters do. To address the issues caused by decoupled parameters, one approach is to employ an adapter that unifies memory across different granularities, another is to fuse the decoupled parameters and dynamically adjust their scaling according to the specific task. Additionally, SAM 2++ still faces challenges in accurately tracking objects under severe occlusion, fast motion, and the presence of similar distractors. To further enhance model performance in these difficult scenarios, introducing motion modeling mechanisms and specialized memory designs could be effective solutions.